# EEG Foundation Challenge: From Cross-Task to Cross-Subject EEG Decoding

### v1

@@ 1 Competition description
In recent years, the neuroscience community has increasingly looked toward Large Language-Model-style architectures to advance brain activity decoding [57]. However, this paradigm shift remains limited by the data scale and structure of currently available data [8, 22]. Unlike the massive, richly annotated corpora used to train language models, EEG datasets are relatively small and lack comparable hierarchical organization. Moreover, brain signal data present intrinsic challenges, including differences in electrode montages, recording protocols, various methods for splitting the data, and underlying neurological factors [6, 11, 17, 31, 51]. These challenges have hindered the development of broad, generalizable, and cross-task models that extend beyond simple classification, highlighting the need for larger, more integrated EEG datasets to support next-generation research.
A given EEG dataset usually corresponds to one specific cognitive task or condition with only a single type of ground-truth label, often requiring a separate model to be trained for each task. In practice, models are frequently developed and tuned in isolation on each dataset—and sometimes even on each individual subject—due to pronounced physiological differences between subjects [3, 36, 44]. This conventional single-task, -subject training and evaluation fails to leverage commonalities across datasets and yields models that struggle to generalize beyond the context in which they were trained.
@@
Previous EEG decoding challenges have shown that well-designed competitions can drive progress in brain decoding. Our BEETL Motor Imagery Challenge at NeurIPS 2021 [54] gathered 130+ contestants from 40 research groups and over 1,382 submissions. Our Brain Age Prediction Challenge 2022 [37] attracted 200+ participants from 40 countries, generating 30–100 daily submissions. In 2023, our Sleep States Competition [26] engaged 80 teams from 20 countries. In 2024, we co-organized a large-scale challenge with the Child Mind Institute [46], focusing on predicting problematic internet use in youth using accelerometry, assessments, and questionnaires. It drew 12,912 registrants, 2,436 active contestants across 1,877 teams, and 38,329 submissions. This year, we are currently running additional challenges on the Healthy Brain Network’s fMRI and actigraphy. The ongoing WiDS Datathon 2025 [10], part of the Women in Data Science initiative, has already reached 5,062 contestants, with 1,614 active contestants across 703 teams and 9,470 submissions to date. These results highlight our strong track record of organizing impactful, large-scale challenges and mobilizing the machine learning research community for an appealing NeurIPS competition.
@@ 1.1 Novelty
1) Unprecedented Scale and Complexity: The competition’s dataset encompasses over 3,000 participants with high-density 128-channel EEG recordings – an order of magnitude larger than typical EEG challenge datasets. Each participant engaged in six distinct cognitive tasks (ranging from resting-state to active learning and attention tasks), providing a rich multi-task, multi-condition collection of neural data. This participant, metadata, and recording breadth and diversity far exceeds that of any prior EEG competition [26, 54, 37]. So, models should handle heterogeneous time-series data conditions, learn generalizable representations across tasks and participants. EEG decoding research has progressed beyond traditional discrete classification methods, with novel approaches now addressing more complex representational challenges at this large scale [57, 8, 7, 53, 52, 32, 16, 23].
2) Zero-Shot Cross-Domain Generalization: The cross-task transfer learning scenario in EEG decoding is remarkably underexplored [5]. Our Challenge 1 addresses a key goal in neurotechnology: decoding cognitive function from EEG without explicit behavioural labels. Participants must develop models capable of zero-shot generalization across both novel tasks and novel subjects. The competition is structured to reward models that learn domain-invariant and subject-invariant representations. This means a submitted model might be trained on a subset of tasks and then tested on data from a held-out task or condition, evaluating its capacity to generalize without task-specific fine-tuning. Notably, no existing benchmarks or datasets have been developed to address these issues [14, 20].
3) Prediction of Latent Psychological Constructs: In contrast to past challenges, our Challenge 2 targets latent psychological constructs as prediction outputs. Specifically, participants will predict transdiagnostic psychopathology factors (such as the p-factor) derived from standardized clinical questionnaires [48, 12]. While the p-factor is intended to reflect a psychological trait [12], EEG could be also modulated by fluctuating “states” such as engagement, fatigue, and other time-variant components [9, 43]. These components may be implicitly encoded in neural activity, helping to account for such variabilities in the data. The challenge holds considerable potential for scientific impact, particularly if participating teams could improve decoding of psychopathology factors from EEG data, thereby offering empirical support for EEG-based biomarkers for mental health [27].
@@ 1.2 Data
The competition leverages the Healthy Brain Network Electroencephalography (HBN-EEG) datasets [49]. It is made of 12 dataset releases in total, grouped according to the time of release, 11 releases shared, and one withheld and unreleased for competitions. This large-scale collection contains high-density (128-channel) EEG recordings from children and young adults aged 5-21 years with more details described at [49]. The dataset is formatted according to the Brain Imaging Data Structure (BIDS) standard [24, 40] and includes comprehensive event annotations using Hierarchical Event Descriptors (HED) [42, 35], making it particularly suitable for cross-task analysis and machine learning applications.
Formally, let $X\in\mathbb{R}^{c\times t}$ be an EEG recording segment (Figure 1), with $c=128$ the number of electrodes and $t$ the time steps, $Y\in\mathbb{R}^{k\times t}$ be the associated labels, with $k$ the number of event labels, and $P\in\mathbb{R}^{7}$ be subject’s traits including 3 demographic attributes and 4 psychological factors, e.g., p-factor $\mathbf{p}_{fac}$. A dataset is defined as $\left({\color[rgb]{0.4453125,0.55859375,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.4453125,0.55859375,0.734375}\mathcal{X}},{\color[rgb]{0.5546875,0.6953125,0.43359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5546875,0.6953125,0.43359375}\mathcal{Y}},{\color[rgb]{0.5078125,0.390625,0.5859375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5078125,0.390625,0.5859375}\mathcal{P}}\right)$, with $N$ EEG recording from $S$ subjects, ${\color[rgb]{0.4453125,0.55859375,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.4453125,0.55859375,0.734375}\mathcal{X}}=\left\{{\color[rgb]{0.4453125,0.55859375,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.4453125,0.55859375,0.734375}\mathbf{X}}_{i}\right\}_{N}$ with their associated labels ${\color[rgb]{0.5546875,0.6953125,0.43359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5546875,0.6953125,0.43359375}\mathcal{Y}}=\left\{{\color[rgb]{0.5546875,0.6953125,0.43359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5546875,0.6953125,0.43359375}Y}_{i}\right\}_{N}$ and psychological factors and demographic attributes ${\color[rgb]{0.5078125,0.390625,0.5859375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5078125,0.390625,0.5859375}\mathcal{P}}=\left\{{\color[rgb]{0.5078125,0.390625,0.5859375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5078125,0.390625,0.5859375}\mathbf{p}}\right\}_{7}$ for each $s$ subject.
We consider here different datasets with different types of labels. In this competition, the trial timing is assumed to be known, targeting offline applications that are common in the medical context, leaving real-time decoders outside the scope of this competition. We assume that the dataset is randomly split into a training set of size $N_{\mathrm{train}}$, a validation set of size $N_{\mathrm{valid}}$ and a test set of size $N_{\mathrm{test}}$. For training and validation sets, both $\mathcal{X}$ and $\mathcal{Y}$ are available, while for test sets, only $\mathcal{X}$ is available to the contestant. The general decoding problem amounts to using the training set ${\color[rgb]{0.4453125,0.55859375,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.4453125,0.55859375,0.734375}\mathcal{X}}_{\mathrm{train}}$ to learn a model $f_{\theta}$ with parameters $\theta$ mapping each trial ${\color[rgb]{0.4453125,0.55859375,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.4453125,0.55859375,0.734375}\mathbf{X}}$ to the associated label ${\color[rgb]{0.5546875,0.6953125,0.43359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5546875,0.6953125,0.43359375}\mathbf{Y}}$.
@@ 1.2.1 Dataset Description
The data of each participant are accompanied by four dimensions of psychopathology (internalization, externalization, attention, and p-factor) derived from a bifactor model of parent-reported questionnaire responses to the Child Behavior Checklist (CBCL) [1, 48]. These psychopathology factors represent orthogonal dimensions of mental health and serve as target variables for the regression component for Challenge 2 [48]. Additionally, the age, gender, and handedness (scored using the Edinburg Inventory, EHQ-Total) [38] are included as part of the demographic information for each participant. While some psychopathology factors, like internalizing and externalizing, are influenced by a participant’s age, other factors such as the p-factor and attention, are largely unaffected by age Figure 1B.
@@
The first 11 releases of the HBN-EEG dataset are freely available under CC-BY-SA-4.0 licence through NEMAR [18]. $\mathrm{Testing}$ set use release 12 that is currently unreleased. The Chesapeake Institutional Review Board approved the data collection, with written informed consent obtained from participants 18 years or older and from legal guardians of younger participants. All participants were anonymized using the NIMH’s Global Unique Identifier Tool (GUID) [30] to ensure no exposure to personally identifiable information.
@@ 1.2.2 Competition Data Split
As contestants submit their code, they will not have direct access to test set. No subject’s data is spread across multiple releases; this means the $\mathrm{Training}$, $\mathrm{Validation}$ and $\mathrm{Testing}$ are completely separate in terms of subjects, ensuring no feature leakage in this competition, avoid subjects co-founders. An illustration of the data split can be found in Figure 1.
For the Challenges, participants are strongly encouraged to leverage additional datasets beyond the provided training set, particularly those involving SSVEP stimuli. As a starting point, we recommend exploring the Mother of All BCI Benchmarks (MOABB) library [14, 4], which provides a wide range of small, but well-curated EEG datasets. However, contestants are responsible to design models capable of adapting to variations in channel configurations and signal characteristics across datasets.
@@ 1.3 Tasks and application scenarios
[…]
In Challenge 1 - Transferability across subjects and cognitive tasks, the contestants are provided with multiple databases with subjects performing similar tasks, namely SSVEP, to train their model on the HBN Releases datasets. For this challenge, the contestants must predict behavioural performance metrics in an active task (Contrast Change Detection, CCD): response time relative to the start of contrast change and success rate (hit accuracy). The test data will consist of test subject EEG data from a passive task (Surround Suppression, SuS), the two-second data epochs preceding each contrast change, plus their demographic and psychopathologic factors, as indicated in Figure 1. Detailed data split, epoching and basic scripts for data loading and curation are available at the comptetion website. Contestants should submit per-trial predictions. The metrics used to evaluate the predictions are detailed in Sect 1.4.
This challenge has significant implications for clinical and research settings as obtaining and retaining EEG recordings and accounting for participant’s response variability may be difficult or impossible due to participants’ or environment limitations [21, 19]. The experiment length and actively engaging participants in experiments are challenging for several populations, including children, individuals with severe motor impairments, or those with communication and developmental disorders [21, 56, 41]. Passive EEG paradigms such as movie watching with shorter duration that can predict performance capabilities would enable more inclusive and accessible neurological assessments. Additionally, by extracting maximal information from minimal recording sessions, researchers can reduce the time and effort required from participants, making neurophysiological research more accessible and efficient.
@@
This challenge represents a novel approach to mental health assessment that leverages objective neurophysiological measures instead of relying on subjective self-reports or clinical observations. It addresses critical challenges in mental health assessment and treatment. Current psychiatric diagnoses mainly rely on subjective assessments, leading to potential inconsistencies and biases [34]. EEG-based biomarkers could provide complementary, physiologically-grounded measures of mental health [9]. Neurophysiological signatures of psychopathology may be detectable before clinical symptoms fully manifest, potentially enabling earlier intervention. By predicting continuous psychopathology dimensions rather than categorical diagnoses, this approach aligns with modern transdiagnostic frameworks like the Research Domain Criteria that conceptualizes mental health along continuous dimensions [28, 15]. Also, objective EEG-based measures could be used as biomarkers to track treatment response over time, providing quantitative feedback on intervention efficacy.
@@ 1.4 Metrics
[…]
@@
Contestants will develop models to predict the psychopathology factors from EEG recordings. The evaluation metric $\mathcal{S}_{2}$ for this challenge is the combination following: the Concordance Correlation Coefficient $\mathcal{S}_{\text{CCC}}$ is the metric measuring both correlation and agreement between predicted and actual psychopathology scores. It penalizes both random scatter and systematic bias. The Root Mean Square Error $\mathcal{S}_{\text{rmse}}$ is the metric quantifying prediction error magnitude, providing a clinically interpretable measure of prediction deviation. Finally, the Spearman’s Rank Correlation $\mathcal{S}_{\text{SRC}}$ is the metric assessing preservation of ranking along each psychopathology dimension. This metric is important for clinical triage applications. The overall weighted metric is:
$\mathcal{S}_{2}=0.5\,\mathcal{S}_{\text{CCC}}+0.3\,\mathcal{S}_{\text{rmse}}+0.2\,\mathcal{S}_{\text{SRC}}$ (2)
@@
The objective of this competition being to foster the development of foundation models, the overall ranking of this competition is a combination of both challenges to promote the design of models able to address multiple tasks. The final metric is: $\mathcal{S}_{\text{overall}}=0.4\,\mathcal{S}_{1}+0.6\,\mathcal{S}_{2}$, reflecting the greater clinical significance of psychopathology prediction. This comprehensive evaluation framework ensures that winning approaches demonstrate robust performance across multiple clinically relevant dimensions, encouraging the development of models with real-world utility for neurophysiological assessment and mental health applications.

### v2

@@ 1 Competition description
In recent years, the neuroscience community has increasingly looked toward Large Language-Model-style architectures to advance brain activity decoding Zhou et al. (2025). However, this paradigm shift remains limited by the data scale and structure of currently available data Bomatter and Gouk (2025); Gemein et al. (2025). Unlike the massive, richly annotated corpora used to train language models, EEG datasets are relatively small and lack comparable hierarchical organization. Moreover, brain signal data present intrinsic challenges, including differences in electrode montages, recording protocols, various methods for splitting the data, and underlying neurological factors Aristimunha et al. (2024); Carlson et al. (2025); Delorme (2023); Kessler et al. (2024); Trübutschek et al. (2024). These challenges have hindered the development of broad, generalizable, and cross-task models that extend beyond simple classification, highlighting the need for larger, more integrated EEG datasets to support next-generation research.
A given EEG dataset usually corresponds to one specific cognitive task or condition with only a single type of ground-truth label, often requiring a separate model to be trained for each task. In practice, models are frequently developed and tuned in isolation on each dataset—and sometimes even on each individual subject—due to pronounced physiological differences between subjects (Apicella et al., 2022; Mikkelsen et al., 2021; Saha and Baumert, 2019). This conventional single-task, -subject training and evaluation fails to leverage commonalities across datasets and yields models that struggle to generalize beyond the context in which they were trained.
@@
Previous EEG decoding challenges have shown that well-designed competitions can drive progress in brain decoding. Our BEETL Motor Imagery Challenge at NeurIPS 2021 (Wei et al., 2022) gathered 130+ contestants from 40 research groups and over 1,382 submissions. Our Brain Age Prediction Challenge 2022 (NeuroTechX and network, 2022) attracted 200+ participants from 40 countries, generating 30–100 daily submissions. In 2023, our Sleep States Competition (Hackathon, 2023) engaged 80 teams from 20 countries. In 2024, we co-organized a large-scale challenge with the Child Mind Institute (Santorelli et al., 2024), focusing on predicting problematic internet use in youth using accelerometry, assessments, and questionnaires. It drew 12,912 registrants, 2,436 active contestants across 1,877 teams, and 38,329 submissions. This year, we are currently running additional challenges on the Healthy Brain Network’s fMRI and actigraphy. The ongoing WiDS Datathon 2025 (by WiDS Worldwide: https://www.widsworldwide.org/, 2025), part of the Women in Data Science initiative, has already reached 5,062 contestants, with 1,614 active contestants across 703 teams and 9,470 submissions to date. These results highlight our strong track record of organizing impactful, large-scale challenges and mobilizing the machine learning research community for an appealing NeurIPS competition.
@@ 1.1 Novelty
1) Unprecedented Scale and Complexity: The competition’s dataset encompasses over 3,000 participants with high-density 128-channel EEG recordings – an order of magnitude larger than typical EEG challenge datasets. Each participant engaged in six distinct cognitive tasks (ranging from resting-state to active learning and attention tasks), providing a rich multi-task, multi-condition collection of neural data. This participant, metadata, and recording breadth and diversity far exceeds that of any prior EEG competition (Hackathon, 2023; Wei et al., 2022; NeuroTechX and network, 2022). So, models should handle heterogeneous time-series data conditions, learn generalizable representations across tasks and participants. EEG decoding research has progressed beyond traditional discrete classification methods, with novel approaches now addressing more complex representational challenges at this large scale Zhou et al. (2025); Bomatter and Gouk (2025); Banville et al. (2025); Wang et al. (2025); Turgut et al. (2025); Kiessner et al. (2024); d’Ascoli et al. (2024); Gemein et al. (2020).
2) Zero-Shot Cross-Domain Generalization: The cross-task transfer learning scenario in EEG decoding is remarkably underexplored Aristimunha et al. (2023b). Our Challenge 1 addresses a key goal in neurotechnology: decoding cognitive function from EEG without explicit behavioural labels. Participants must develop models capable of zero-shot generalization across both novel tasks and novel subjects. The competition is structured to reward models that learn domain-invariant and subject-invariant representations. This means a submitted model might be trained on a subset of tasks and then tested on data from a held-out task or condition, evaluating its capacity to generalize without task-specific fine-tuning. Notably, no existing benchmarks or datasets have been developed to address these issues Chevallier et al. (2024); Ding et al. (2021).
3) Prediction of Latent Psychological Constructs: In contrast to past challenges, our Challenge 2 targets latent psychological constructs as prediction outputs. Specifically, participants will predict transdiagnostic psychopathology factors (such as the p-factor) derived from standardized clinical questionnaires (Scopel Hoffmann et al., 2022; Caspi et al., 2014). While the p-factor is intended to reflect a psychological trait Caspi et al. (2014), EEG could also be modulated by fluctuating “states” such as engagement, fatigue, and other time-variant components Bomatter et al. (2024); Sabbagh et al. (2020). These components may be implicitly encoded in neural activity, helping to account for such variabilities in the data. The challenge holds considerable potential for scientific impact, particularly if participating teams could improve decoding of psychopathology factors from EEG data, thereby offering empirical support for EEG-based biomarkers for mental health Hong et al. (2023).
@@ 1.2 Data
The competition leverages the Healthy Brain Network Electroencephalography (HBN-EEG) datasets Shirazi et al. (2024). It is made of 12 dataset releases in total, grouped according to the time of release, 11 releases shared, and one withheld and unreleased for competitions. This large-scale collection contains high-density (128-channel) EEG recordings from children and young adults aged 5-21 years with more details described at (Shirazi et al., 2024). The dataset is formatted according to the Brain Imaging Data Structure (BIDS) standard (Gorgolewski et al., 2016; Pernet et al., 2019) and includes comprehensive event annotations using Hierarchical Event Descriptors (HED) (Robbins et al., 2021; Makeig and Robbins, 2024), making it particularly suitable for cross-task analysis and machine learning applications.
[…]
@@ 1.2.1 Dataset Description
The data of each participant are accompanied by four dimensions of psychopathology (internalization, externalization, attention, and p-factor) derived from a bifactor model of parent-reported questionnaire responses to the Child Behavior Checklist (CBCL) (Achenbach, 1999; Scopel Hoffmann et al., 2022). These psychopathology factors represent orthogonal dimensions of mental health and serve as target variables for the regression component for Challenge 2 Scopel Hoffmann et al. (2022). Additionally, the age, gender, and handedness (scored using the Edinburg Inventory, EHQ-Total) (Oldfield, 1971) are included as part of the demographic information for each participant. While some psychopathology factors, like internalizing and externalizing, are influenced by a participant’s age, other factors such as the p-factor and attention, are largely unaffected by age Figure 1B.
@@
The first 11 releases of the HBN-EEG dataset are freely available under CC-BY-SA-4.0 licence through NEMAR (Delorme et al., 2022). $\mathrm{Testing}$ set use release 12 that is currently unreleased. The Chesapeake Institutional Review Board approved the data collection, with written informed consent obtained from participants 18 years or older and from legal guardians of younger participants. All participants were anonymized using the NIMH’s Global Unique Identifier Tool (GUID) (Johnson et al., 2010) to ensure no exposure to personally identifiable information.
@@ 1.2.2 Competition Data Split
As contestants submit their code, they will not have direct access to the test set. No subject’s data is spread across multiple releases; this means the $\mathrm{Training}$, $\mathrm{Validation}$, and $\mathrm{Testing}$ are completely separate in terms of subjects, ensuring no feature leakage in this competition, avoiding subjects’ co-founders. An illustration of the data split can be found in Figure 1.
For the Challenges, participants are strongly encouraged to leverage additional datasets beyond the provided training set, particularly those involving SSVEP stimuli. As a starting point, we recommend exploring the Mother of All BCI Benchmarks (MOABB) library Chevallier et al. (2024); Aristimunha et al. (2023a), which provides a wide range of small, but well-curated EEG datasets. However, contestants are responsible for designing models capable of adapting to variations in channel configurations and signal characteristics across datasets.
@@ 1.3 Tasks and application scenarios
$\text{Challenge 1 - }f_{\theta}:\left({\color[rgb]{0.4453125,0.55859375,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.4453125,0.55859375,0.734375}\mathcal{X}},{\color[rgb]{0.5078125,0.390625,0.5859375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5078125,0.390625,0.5859375}\mathcal{P}}\right)\to{\color[rgb]{0.5546875,0.6953125,0.43359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5546875,0.6953125,0.43359375}\mathcal{Y}},\;\;\;\text{Challenge 2 - }f_{\theta}:\left({\color[rgb]{0.4453125,0.55859375,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0.4453125,0.55859375,0.734375}\mathcal{X}},{\color[rgb]{0.5546875,0.6953125,0.43359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5546875,0.6953125,0.43359375}\mathcal{Y}}\right)\to\ {\color[rgb]{0.5078125,0.390625,0.5859375}\definecolor[named]{pgfstrokecolor}{rgb}{0.5078125,0.390625,0.5859375}\mathcal{P}},$
In Challenge 1 - Transferability across subjects and cognitive tasks, the contestants are provided with multiple databases with subjects performing similar tasks, namely SSVEP with ERP components, to train their model on the HBN Releases datasets. For this challenge, the contestants must predict behavioural performance metrics in an active task (Contrast Change Detection, CCD): response time relative to the start of contrast change index on the stimulus time with a 500 ms shift, i.e., The window period is 0.5 to 2.5 seconds after stimulus onset for prediction.
At the inference time, we will use the Contrast Change Detection data with the two-second data epochs and plus their demographic at the epoch level, as indicated in Figure 1.
Detailed data split, epoching, and basic scripts for data loading and curation are available at the competition website. Contestants should submit per-trial predictions. The metrics used to evaluate the predictions are detailed in Sect 1.4.
This challenge has significant implications for clinical and research settings as obtaining and retaining EEG recordings and accounting for participant’s response variability may be difficult or impossible due to participants’ or environment limitations (DiStefano et al., 2019; Demazure et al., 2021). The experiment length and actively engaging participants in experiments are challenging for several populations, including children, individuals with severe motor impairments, or those with communication and developmental disorders (DiStefano et al., 2019; Zech et al., 2022; Reilly et al., 2019). Passive EEG paradigms such as movie watching with shorter duration that can predict performance capabilities would enable more inclusive and accessible neurological assessments. Additionally, by extracting maximal information from minimal recording sessions, researchers can reduce the time and effort required from participants, making neurophysiological research more accessible and efficient.
@@
This challenge represents a novel approach to mental health assessment that leverages objective neurophysiological measures instead of relying on subjective self-reports or clinical observations. It addresses critical challenges in mental health assessment and treatment. Current psychiatric diagnoses mainly rely on subjective assessments, leading to potential inconsistencies and biases Loo et al. (2016). EEG-based biomarkers could provide complementary, physiologically-grounded measures of mental health Bomatter et al. (2024). Neurophysiological signatures of psychopathology may be detectable before clinical symptoms fully manifest, potentially enabling earlier intervention. By predicting continuous psychopathology dimensions rather than categorical diagnoses, this approach aligns with modern transdiagnostic frameworks like the Research Domain Criteria that conceptualizes mental health along continuous dimensions Insel et al. (2010); Cuthbert (2015). Also, objective EEG-based measures could be used as biomarkers to track treatment response over time, providing quantitative feedback on intervention efficacy.
@@ 1.4 Metrics
Contestants will predict behavioural performance, i.e., the response time, in the Contrast Change Detection (CCD) task using EEG data from the Surround Suppression (SuS) task and pre-trial EEG. We decided not to use correct/incorrect classifications for this challenge. The evaluation metric for this regression challenge is the normalized root mean square error $\text{nRMSE}\left(\cdot,\cdot\right)$ between the true and predicted response time. The normalized root mean square error is the metric quantifying prediction error magnitude, providing a clinically interpretable measure of prediction deviation Formally, this metric is defined as follows:
$\text{RMSE}\left(\mathbf{Y}_{\text{true}},\mathbf{Y}_{\text{pred}}\right)$ $=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left|\mathbf{Y}^{i}_{\text{true}}-\mathbf{Y}^{i}_{\text{pred}}\right|^{2}}$ (1)
$\text{std}\left(\mathbf{Y}\right)$ $=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left|\mathbf{Y}^{i}-\overline{\mathbf{Y}}\right|^{2}}$ (2)
$\text{nRMSE}\left(\mathbf{Y}_{\text{true}},\mathbf{Y}_{\text{pred}}\right)$ $=\frac{\text{RMSE}\left(\mathbf{Y}_{\text{true}},\mathbf{Y}_{\text{pred}}\right)}{\text{std}\left(\mathbf{Y}_{\text{true}}\right)}$ (3)
$\mathcal{S}_{1}$ $=\text{nRMSE}\left(\mathbf{Y}_{1,\text{true}},\mathbf{Y}_{1,\text{pred}}\right)$ (4)
With $\mathbf{Y}_{1,\text{true}}$ being the vector containing the true response times, $\mathbf{Y}_{1,\text{pred}}$ the predicted ones, and $\mathcal{S}_{1}$ the score for challenge 1.
@@
Contestants will develop models to predict the psychopathology factors from EEG recordings. The evaluation metric for this challenge is also the normalized root mean square error:
$\mathcal{S}_{2}=\text{nRMSE}\left(\mathbf{Y}_{2,\text{true}},\mathbf{Y}_{2,\text{pred}}\right)$ (5)
With $\mathbf{Y}_{2,\text{true}}$ being the labels vector containing the p-factors obtained from questionaries, $\mathbf{Y}_{2,\text{pred}}$ containing the p-factors predicted from the EEG recordings, and $\mathcal{S}_{2}$ the score for challenge 2.
@@
The objective of this competition being to foster the development of foundation models, the overall ranking of this competition is a combination of both challenges to promote the design of models able to address multiple tasks. The final score $\mathcal{S}_{\text{overall}}$ reflects the greater clinical significance of psychopathology prediction. This comprehensive evaluation framework ensures that winning approaches demonstrate robust performance across multiple clinically relevant dimensions, encouraging the development of models with real-world utility for neurophysiological assessment and mental health applications.
$\mathcal{S}_{\text{overall}}=0.3\,\mathcal{S}_{1}+0.7\,\mathcal{S}_{2}$ (6)